OpenStudy (anonymous):

How would you solve \[4ln(x) - 3 ln(2x)= 10\]?

OpenStudy (anonymous):

PLZ SOLVE THS QUESTION

OpenStudy (anonymous):

\[\ln x=10+3\ln 2 \] \[x=8e^{10}\]

OpenStudy (anonymous):

ln(x^4)-ln(8*x^3)=10 ln(x/8)=10 x/8=e^10 x=8*e^10

OpenStudy (valpey):

Is the equation editor broken? 4ln(x) - 3 ln(2x) = 10 ln(x^4) - ln((2x)^3) = 10 ln((x^4)/(8x^3)) = 10 ln(x/8) = 10 x/8 = e^10 x = 8e^10

OpenStudy (anonymous):

Oh that makes sense...I feel dumb :P Thanks

OpenStudy (anonymous):

i mentally solved it ...u see thing is : 3ln(2x)=3{ln(2)+ln(x)} now wen this is subtracted from 4ln(x) we get ln(x)-3*ln(2)=10 shifting 3*ln(2) on rhs.... \[x=e^{10}\]